# Li Tana,b , Shengrong Wanga,c

### v1

@@ 5.1 Moment bounds
Let Assumption 2.2-2.5 and Lemmas 5.1, 5.2 hold. For $\overline{p}>1/H$, we have
@@
For $a>0$, let $\lfloor a\rfloor$ be the integer part of $a$. Applying the Itô formula to $[1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$, we get
@@
$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+p\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
$+\overline{p}(\overline{p}-2)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\sigma(\overline{\mu}^{Y,N}(s))D_{s}^{\phi}[1+|\widetilde{Y}_{\Delta}(s)|^{2}]{\rm d}s$
$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+\overline{p}(\overline{p}-2)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|\sigma(\overline{\mu}^{Y,N}(s))|^{2}\int_{0}^{s}\phi(s,r){\rm d}r{\rm d}s$
@@
=:absent:=:= : $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+G_{1}(t)+G_{2}(t)+G_{3}(t).$
where $\widehat{Y}_{\Delta}^{i,N}(s)=\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\Delta$. By the properties of $\phi$ and the Young’s inequality, Lemma 5.1, we get
$G_{1}(t)$
$\leq$ $CHT^{2H-1}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|\sigma(\overline{\mu}^{Y,N}(s))|^{2}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\big{(}1+[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}}{2}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\big{)}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg{)}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg{)}{\rm d}s$
@@
By Remark 5.3 and the Young’s inequality, we get
$G_{2}(t)\leq$ $C\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{(}1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2}+(\mathbb{W}_{q}(\mu^{\overline{Y},N}(s),\delta_{0}))^{2}\bigg{)}{\rm d}s$
@@
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
=:absent:=:= : $\overline{p}G_{31}(t)+\overline{p}G_{32}(t),$
where $\widetilde{Y}_{\Delta}^{i,N}(t)-\widehat{Y}_{\Delta}^{i,N}(s)=\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u+\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}$. By Lemma 5.1 and the Young’s inequality, we get
@@
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|{\rm d}s$
@@
Applying Itô formula to $[1+|\widetilde{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$, we get
@@
$+(\overline{p}-2)(\overline{p}-4)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{r}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
@@
Then,
$[1+|\widehat{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
@@
$+(\overline{p}-2)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
$+(\overline{p}-2)(\overline{p}-4)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{u}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
$+(\overline{p}-2)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
@@
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$+(\overline{p}-2)(\overline{p}-4)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{u}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
@@
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
=:absent:=:= : $(\overline{p}-2)G_{321}(t)+(\overline{p}-2)(\overline{p}-4)G_{322}(t)+(\overline{p}-2)G_{323}(t).$
@@
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
@@
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$\leq$ $\Delta\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{3}{\rm d}u{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u),\overline{\mu}^{Y,N}(u))|{\rm d}u$
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|\bigg{)}^{\frac{\overline{p}}{\overline{p}-1}}+\bigg{\lvert}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}\bigg{\rvert}^{\overline{p}}\bigg{]}{\rm d}s$
@@
$+C\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}u|\bigg{)}^{\frac{\overline{p}}{\overline{p}-1}}$
@@
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
Similarly, we get
$G_{322}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
By Assumption 2.4, Lemma 5.1 and [32, Lemma 3.1.3], noting that $\mathbb{E}\int_{0}^{t}F_{s}{\rm d}B_{s}^{H,i}=0$, we get
$G_{323}(t)=$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$

### v2

@@ 5.1 Moment Boundness
Let Assumptions 2.2-2.5 hold. For $\overline{p}\geq 2$, we have
@@
We begin by considering the case $\overline{p}\geq 4$, using moment estimation techniques for numerical solutions from [36]. The case $2\leq\overline{p}<4$ then follows directly from the Young’s inequality. For $a>0$, let $\lfloor a\rfloor$ be the integer part of $a$. Applying the fractional Itô formula to $1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$ and using the fact that $\mathbb{E}B_{s}^{H,i}=0$, it holds that
@@
$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
$+\overline{p}(\overline{p}-2)H(2H-1)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(r))\|^{2}{\rm d}r\bigg){\rm d}s$
$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+\overline{p}(\overline{p}-2)H(2H-1)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
$\times\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(r))\|^{2}{\rm d}r\bigg){\rm d}s$
@@
$=:$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+G_{1}(t)+G_{2}(t)+G_{3}(t).$
where
$\widehat{Y}_{\Delta}^{i,N}(s)=\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\Delta.$
For $t\in[0,T]$, the Young’s inequality, the Fubini theorem and Lemma 5.1 give that
$G_{1}(t)\leq$ $C\mathbb{E}\int_{0}^{t}\bigg(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}$
$+|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta|^{\overline{p}}\bigg){\rm d}s+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}|\overline{Y}_{\Delta}^{i,N}(r)|^{\overline{p}}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg){\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s$
@@
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
Remark 5.3 together with the Young’s inequality yield
$G_{2}(t)\leq$ $C\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2}+\mathbb{W}_{2}^{2}(\mu^{\overline{Y},N}(s),\delta_{0})\bigg){\rm d}s$
@@
$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$=:$ $\overline{p}G_{31}(t)+\overline{p}G_{32}(t),$
where
$\widetilde{Y}_{\Delta}^{i,N}(t)-\widehat{Y}_{\Delta}^{i,N}(s)=\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u+\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}.$
We use Lemma 5.1 and the Young’s inequality to get
@@
$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|{\rm d}s$
@@
Again, by the fractional Itô formula
@@
$+\frac{1}{2}H(2H-1)(\overline{p}-2)(\overline{p}-3)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(r)|^{2}]^{\frac{\overline{p}-4}{2}}\int_{0}^{r}(r-u)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(u))\|^{2}{\rm d}u{\rm d}r$
@@
Then, by the Young’s inequality and the Fubini theorem, we have
$\mathbb{E}[1+|\widehat{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
@@
$+(\overline{p}-2)\mathbb{E}\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
$+\frac{1}{2}H(2H-1)(\overline{p}-2)(\overline{p}-3)\bigg(\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}\frac{u^{2H-1}}{2H-1}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}}{2}}{\rm d}u$
$+\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}\frac{(\lfloor\frac{s}{\Delta}\rfloor\Delta-u)^{2H-1}}{2H-1}\mathbb{E}|\overline{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}u\bigg)$
$+(\overline{p}-2)\mathbb{E}\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
@@
$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$+(\overline{p}-2)(\overline{p}-3)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\bigg(\frac{u^{2H-1}}{2H-1}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}}{2}}$
$+\frac{(s-u)^{2H-1}-(\lfloor\frac{s}{\Delta}\rfloor\Delta-u)^{2H-1}}{2H-1}\mathbb{E}|\overline{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}u$
$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
@@
$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$=:$ $(\overline{p}-2)G_{321}(t)+(\overline{p}-2)(\overline{p}-3)G_{322}(t)+(\overline{p}-2)G_{323}(t).$
@@
$\times\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle{\rm d}s$
@@
$\times\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle{\rm d}s$
$\leq$ $\Delta\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{3}{\rm d}u{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\bigg[\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))|{\rm d}u$
$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|\bigg)^{\frac{\overline{p}}{\overline{p}-1}}+\bigg\lvert\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}\bigg\rvert^{\overline{p}}\bigg]{\rm d}s$
@@
$+C\mathbb{E}\int_{0}^{t}\bigg[\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}u|\bigg)^{\frac{\overline{p}}{\overline{p}-1}}$
@@
Similarly, $G_{322}(t)$ can be estimated by Lemma 2.8 as
$G_{322}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
Using Assumption 2.4, Lemma 5.1 and [28, Lemma 3.1.3], Hölder inequality, Young’s inequality, we derive
$G_{323}(t)$
$\times\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle{\rm d}s$